The Search for the Top Quark

I Introduction II Direct Searches at e^{+} e^{−} Colliders III The CDF Top Quark Search IV Results from the CERN p p Collider V Future Prospects VI Conclusions

Introduction

The standard model has been extraordinarily successful

• Charged and neutral current ν scattering

• Parity violation in heavy nuclei

• Properties of W ± and Z∘ bosons Masses and widths Production and decay properties

• Forward-backward asymmetries in e+ e− → ff

However, it is still a phenomenological theory

• Too many parameters (∼ 21)

• Unification of forces is incomplete

• Higgs mechanism not well understood

• “Heirarchy problem”

The top quark and Higgs have not been observed The standard model requires a top quark to account for many phenomena:

• asymmetry in e+ e− → b b scattering

• limits on FCNC in b decays

• B∘ − B̄∘ mixing

• The measured masses of the W ± and Z∘ and electroweak

radiative corrections require a top quark with mass Mtop ∼ 130 ± 35 GeV/c2

Top Search Techniques

Three ways of creating top quarks

1. Pair production in e+ e− → tt

2. W ± production with W → tb

3. Pair production in pp → tt

The standard model top quark decays:

• semi-leptonically:

t → W + b
|_{⟶} \{ \begin{matrix}e+ νe \\ μ+ νμ \\ τ+ ντ \end{matrix}

• hadronically:

t → W + b
|_{⟶} \{ \begin{matrix}cs̄ \\ ud̄ \end{matrix}

Other modes may be possible in non-minimal models:

• E.g., if a charged Higgs exists, then also could have

t → H+ b
|_{⟶} \{ \begin{matrix}cs̄ \\ τ+ ντ \end{matrix}

Top events are characterized by an unusual number of energetic partons and are very spherical.

• At e^{+} e^{−} colliders, the spherical topology is sufficient to tag top events.

• At hadron colliders, it is necessary to search for specific decay modes to avoid the backgrounds from other QCD processes.

– One t decaying semi-leptonically

– Both t’s decaying semi-leptonically

Searches at e^{+} e^{−} Colliders

Experiments at Tristan, SLC and LEP have reported results of searches for top quark production

• searched for top using various techniques

Total cross section

Event shapes (e.g., aplanarity, thrust)

Isolated high-PT particles

These searches are all limited by the center of mass energy of the collider.

• However, they are insensitive to the detailed decay topologies of the top quark.

⇒ limits hold if a light, charged Higgs exists.

Searches at Tristan have used a combination of

R = \frac{σ(e+ e− → hadrons )}{σ(e+ e− → μ+ μ−)}

• isolated charged tracks

• event shapes

• They have excluded:

Mtop < 29.9 GeV/c2  (Amy )
Mtop < 29.9 GeV/c2  (Topaz )
Mtop < 30.2 GeV/c2  (Venus )

PRL 63, 1910 (1989) PL 234B, 382 (1990) PL 234B, 525 (1990)

Mark II search focused on isolated charged tracks:

• Selected spherical events (Thrust< 0.9)

• Searched for isolated charged tracks

• Observed 1 event (expected 0.9-1.8 from background)

⇒ excludes Mtop < 40.7 GeV/c2 (95%CL) PRL 63, 2447 (1989)

OPAL search focused on event shapes

• Used acoplanarity (> 0.25) to identify top candidates • Backgrounds from Z → qq̄

• Observed 15 events (expected 20-30 background)

⇒ excludes 23.4 < Mtop < 44.5 GeV/c2 (95%CL) PL 236B, 364 (1990)

ALEPH search also used an “isolated track” signature:

• Selected events with acoplanarity> 0.03

• At least one charged particle well-separated from jets

• Observed 6 events and expected 6.5 from background

⇒ excludes 26.0 < Mtop < 45.8 GeV/c2 (95%CL) PL 236B, 511, 1990

DELPHI collaboration considered several different modes

• Selected spherical events (Thrust < 0.9 and “minor”> 0.2)

• Other selections more sensitive to b′ and H+ decay modes

• Observed 48 events and expected 59 ± 10 background

⇒ excludes Mtop < 44.5 GeV/c2 (95%CL) PL 242B, 536, 1990

Summary of e^{+} e^{−} top quark searches

• All the searches are insensitive to the detailed top decay modes

• Exclude a top quark with:

Mtop < 29.9 GeV/c2  (Topaz )
Mtop < 29.9 GeV/c2  (Amy )
Mtop < 30.2 GeV/c2  (Venus )
Mtop < 40.7 GeV/c2  (Mark II)
Mtop < 44.5 GeV/c2  (OPAL )
Mtop < 45.8 GeV/c2  (Aleph )
Mtop < 44.5 GeV/c2  (Delphi )

All of these searches are sensitive to a b′

• 4rth-generation quark lighter than top

• Possible decays via NC diagrams

Most searches have used methods insensitive to b′ BR’s

• Exclude a b′ quark with a mass

Mb′ < 28.4 GeV/c2  (Topaz )
Mb′ < 27.2 GeV/c2  (Amy )
Mb′ < 28.2 GeV/c2  (Venus )
Mb′ < 42.0 GeV/c2  (MarkII)
Mb′ < 45.2 GeV/c2  (OPAL )
Mb′ < 46.0 GeV/c2  (Aleph )
Mb′ < 45.0 GeV/c2  (Delphi )

Top Quark Production in pp Collisions

• t t production cross section has been calculated to order α_{s}^{3} by Nason et al. (NP B303, 607 (1988)).

• Cross sections derived by Altarelli et al. have become an “industry standard”

• uncertainties are of order ±30%. (NP B308, 724 (1988))

• For Mtop > ∼ 70 GeV/c2, W → tb rate is neglible.

The CDF Top Quark Search

The CDF collaboration has performed a search for the top quark in data acquired in 1988-89

• 4.4 pb^{−1} of data

• Most sensitive search technique includes final states with

– lepton + jets

– electron + muon

– e^{+} e^{−} and μ^{+} μ^{−}

The CDF Detector

• The CDF detector is a general purpose, solenoidal detector

• Solenoid magnet of radius ∼ 1.5 m and length 5 m generating a 1.411 T magnetic field.

• Tracking chambers inside magnetic field provide particle

momentum resolution of

\frac{σp}{p} ∼ 0.0011p

• Calorimeters surround the interaction point

• Used for electron and muon identification

• Detection of jets

• Measurement of underlying event

• Detectors outside of the calorimeter used to identify Muons

Lepton, Neutrino + Jets Analysis

• For CDF, the t t final state with an electron, neutrino, and jets is the most sensitive:

– The e± and νe requirement reject most of the QCD multijet background

– 20% of t t events decay into this mode

– The e± and νe identification is one of CDF’s strongest features.

CDF’s first search focused on this mode:

– Characterized the backgrounds

p̄p → W ±X
p̄p → bX
p̄p → jets

– Used sample to understand how backgrounds changed as subsequent neutrino and jets were applied.

Inclusive Electron Selection

• E{em \atop T}> 15 GeV

• E{had \atop T}/E{em \atop T} < 0.05

• lateral shower profile consistent with electron

• strip chamber showers consistent with electron χ2X< 10

χ2Z< 10

• position match with CTC track Δ Rϕ< 1.5 cm Δ Z< 3.0 cm

• momentum match with CTC track ET /PT < 1.5

• fiducial cuts

This resulted in a sample of 17 500 events.

• Selection efficiency measured to be 77 ± 3% for electrons from W and Z∘ decays.

The backgrounds to t t events in this sample come from

1. Electrons from conversions and π∘ → γe+ e−

2. Fake electron candidates from QCD jets

3. Electrons from b and c quark decay

4. Electrons from Z∘ bosons

5. Electrons from W bosons

To remove conversion electrons and electrons from Dalitz decays:

• Removed candidates with a second charged track with m(e+ e−) < 0.5 GeV/c2.

• Removed candidates with no charged track in the innermost tracking chamber.

This selection

• ∼ 95% efficient for prompt electrons

• left a residual conversion background of ∼ 5%.

• Events with electrons from Z^{∘} decays were identified by searching for a 2nd electron candidate.

– Rejected event if m(e+ e−) > 70 GeV/c2.

– These cuts leave an event sample of ∼11 000 events.

• After all the electron cuts, background arising from QCD fakes is estimated to be ∼ 15 − 20%.

This leaves only heavy quarks and W bosons as potential sources of background.

• These have very different electron and neutrino signatures: – Electrons from W ’s are more energetic

– Electrons from semileptonic b decay poorly isolated – W ’s have large E̸T signature

t t candidates were selected by requiring

• 2 or more jets with ET > 10 GeV with η < 2.0.

• Electron E{em \atop T}> 20 GeV and E̸T > 20 GeV

• Results in a sample of 104 events

To quantify the source of these events, used the transverse mass

M_{T}^{eν} ≡ √2E_{T}^{em} E̸_{T} (1 − cos ϕ)

To check understanding of the M_{T}^{eν} variable, considered electron + 1 jet events

• < 20% of this sample could come from top quark production • Has to be described by W +1 jet production.

• The predicted M{eν \atop T} and rate of events agrees well data

Set a statistical limit on the amount of t t production allowed by the data

• performed a likelihood fit of the M{eν \atop T} distribution to the form

\frac{dN}{dM{eν \atop T}} = αT (M{eν \atop T}) + βW (M{eν \atop T})

• Used the likelihood function from the fit to determine the 95% CL upper limit on the rate of produced t t events.

Included various sources of systematic uncertainties on the t t acceptance

1. Integrated luminosity

2. Model of the top quark fragmentation into a meson

3. Uncertainties in the model for t t production

4. Uncertainies in detector resolution and jet energy calibration

• Systematic uncertainties were convoluted together with the observed likelihood distribution from the fit.

• The resulting probability distribution was used to determine the upper limit on the top quark production cross section.

• The upper limits on the t t cross section vary as a function of top quark mass.

• Compared with the theoretical lower limit on the t t cross section as determined by Altarelli et al..

• Exclude at 95% CL a top quark with mass

40 < Mtop < 77 GeV/c2

PRL 64, 192 (1990) PR D43, 664 (1991)

Electron + Muon Analysis

Inclusive Electron Selection

• E{em \atop T}> 15 GeV

• E{had \atop T}/E{em \atop T} < 0.05

• lateral shower profile consistent with electron

• strip chamber showers consistent with electron χ2X< 10

χ2Z< 10

• position match with CTC track Δ Rϕ< 1.5 cm Δ Z< 3.0 cm

• momentum match with CTC track ET /PT < 1.5

• fiducial cuts

Muon Selection

μ candidates selected by requiring a minimum-ionizing charged track:

• CTC track pointing at a calorimeter tower with

Ehad < 6 GeV Eem < 2 GeV Eem + Ehad > 0 GeV

Also required either:

1. track segment in central μ chambers with

• an r − ϕ match with the CTC track of less than 10 cm • CTC track PT > 5 GeV/c

2. or an isolated, high-P_{T} CTC track:

• < 5 GeV energy in calorimeter towers within a cone of

radius R ≡ √Δϕ2 + Δη2 < 0.4 • CTC track PT > 10 GeV/c

Selection efficiency measured using Z^{∘} → μ^{+} μ^{−} events:

0.98 ± 0.02 for μ candidates with μ track stub 0.96 ± 0.02 for μ candidates without μ track stub

One candidate with E_{T}^{e} > 15 GeV and P_{T}^{μ} > 15 GeV.

• Dominant backgrounds in this channel are:

– Z∘ → τ+ τ− → e±μ∓X estimated to contribute ∼ 1 event.

– bottom quark production is estimated to be < 1 event

– W W and W Z production of order 0.1-0.2 events.

• To set a conservative upper limit, the one event is attributed to top production.

• Systematic uncertainties on acceptance are:

– top quark fragmentation Varied ϵ parameter by order of magnitude. 30% uncertainty for 28 GeV/c2 top mass 10% uncertainty for 70 GeV/c2 top mass

– top quark production Compared several different QCD calculations 30% uncertainty for 28 GeV/c2 top mass 4% uncertainty for 70 GeV/c2 top mass

– 15% uncertainty in integrated luminosity

– Added these uncertainties in quadrature and folded them together with a Poisson probability distribution.

• The resulting probability distribution used to determine the 95% CL limits for top quark production.

28 < M_{top} < 72 GeV/c^{2} (95% CL)

PRL 63, 197 (1990)

Combining Limits

CDF has recently extended the mass limit:

• Included lepton + ≥ 2 jet channel with a b → μX tag to reduce W + jet background

• Added the e+ e− and μ+ μ− channels to the eμ analysis • Performed a combined-channel analysis to place a stronger

limit on the t t production cross section

For lepton + jets data, also require that the event have a μ candidate with P_{T} > 2 GeV

• Reduces acceptance by ×5 due to BR(b → μX)

• Additional reduction due to μ acceptance

• Added μ+ ≥ 2 jets to data sample to boost sensitivity by ∼ ×2

• Additional systematic uncertainties are:

– BR(b → μX) (±10%)

– Low energy μ acceptance (±5%)

– Acceptance for 85 GeV t t is 0.228 ± .057%.

The e^{+} e^{−} and μ^{+} μ^{−} cuts were:

• Di-lepton pair mass at least 15 GeV away from MZ

• Require E̸T > 20 GeV

• Azimuthal angle between leptons Δϕ to satisfy 20∘ < Δϕ < 160∘

• Increases the dilepton acceptance by 70%

Only 1 eμ event survives these cuts.

• The dileptons and lepton+jets channels are combined to-

gether to estimate the observed t t cross section:

σt t = \frac{nobs}{∑ ϵi}

• The Poisson fluctuations are folded together with the systematic uncertainies to get an upper limit on the t t cross section.

28 < M_{top} < 91 GeV/c^{2} (95% CL )

Fermilab-PUB-91-280-E (1991)

Limit from Width of the W

W and Z cross section ratio

\frac{σ(W → eν_{e})}{σ(Z → ee)} = \frac{σ(pp̄ → W X)}{σ(pp̄ → Z)} \frac{Γ(W → eν)}{Γ(Z → ee)} \frac{Γ(W)}{Γ(Z)}

• provides a measure of Γ(W)

• places indirect limit on top quark

Selected sub-sample of W and Z events

• 1828 W → eνe and 193 Z → e+ e− events

• \frac{σ(W →eνe)}{σ(Z→ee)} = 10.2 ± 0.8 ± 0.4

• Using Γ(Z) = 2.496 ± 0.016 from LEP, obtain

Γ(W) = 2.20 ± 0.20 GeV

If assume N^{ν} = 3, then exclude M_{top} < 43 GeV/c^{2} (95% CL )

PR D44, 29 (1991)

Searches at the CERN p p Collider

UA2 Collaboration have searched for the top quark decaying into an electron, ν_{e} and several jets.

• On the basis of a similar transverse mass analysis, they exclude a top quark with mass

30 < M_{top} < 69 GeV/c^{2} (95% CL)

Z.Phys. C46, 179 (1990)

UA1 collaboration has searched for the top quark in both

• muon + νμ + jets channel

• the μ+ μ− final state

• They search for both the W → tb and t t final states, and exclude a top quark with

M_{top} < 60 GeV/c^{2} (95% CL )

Z.Phys. C48, 1 (1990)

Future Prospects

• The next round of experiments will take place at the Fermilab Tevatron in 1992

CDF and D0 expect to record ∼ 50 − 100 pb−1

Will extend the search up to Mtop∼ 150 GeV/c2.

• Future Tevatron experiments will be able to extend the search up to a M_{top} ∼ 180 GeV/c^{2}(∼ 1995).

• LEP II will extend limit up to 80 − 90 GeV/c^{2}

• If the top quark is more elusive

– will have to wait for the SSC or LHC to find it.

– However, standard model tests will be precise enough to constrain the top quark mass to within < ∼ 10 GeV/c2.

Conclusions

1. Searches using the lepton + neutrino + jets channels finds no evidence for top production.

40 < Mtop < 77 GeV/c2 (CDF )
30 < Mtop < 69 GeV/c2 (UA2 )
Mtop < 61 GeV/c2 (UA1 )

2. Study of electron + muon channel finds one event.

28 < Mtop < 72 GeV/c2 (CDF)

3. The CDF lepton+jets and dilepton analysis excludes

28 < Mtop < 91 GeV/c2 (CDF)

4. Top searches at the e^{+} e^{−} colliders exclude

Mtop < 29.9 GeV/c2  (Topaz )
Mtop < 29.9 GeV/c2  (Amy )
Mtop < 30.2 GeV/c2  (Venus )
Mtop < 40.7 GeV/c2  (Mark II)
Mtop < 44.5 GeV/c2  (OPAL )
Mtop < 45.8 GeV/c2  (Aleph )
Mtop < 44.5 GeV/c2  (Delphi )

5. The 1992-3 Tevatron run will extend the search up to ∼ 150 GeV/c^{2}.